OpenStudy (anonymous):

What is the limit of (1/t - 1/t^2+t) as x approaches 0?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{1}{t} - \frac{1}{t^2 + t}\]

OpenStudy (anonymous):

I would first try to find a common denominator

OpenStudy (zzr0ck3r):

I assume you mean when t goes to 0?

OpenStudy (anonymous):

Yea, that's what the question asks, right? Or is it asking for x -> 0?

OpenStudy (anonymous):

\[\frac{1}{t}-\frac{1}{t+t^2}=\frac{1}{1+t} \]